Overproduction of uric acid in the body can be an indication of cell breakdown. This may be an advanced indication of illness such as gout, leukemia, or lymphoma. Over a period of months, an adult male patient has taken eight blood tests for uric acid. The sample mean concentration was 5.33 mg/dL. The distribution of uric acid in healthy adult males can be assumed to be normal, with population standard deviation 1.85 mg/dL.

In steps we are going to find a 95% confidence interval for the population mean.

1) In order to find any confidence interval in the chapter, you must calculate the EBM (error bound) value. In this case you must identify the critical value to be used for ZCL

ZCL  = ____________

2) What is the error bound for a population mean, EBM, for this problem (rounded to the nearest hundredth)?

____________________
Answer #1

a)

sample mean, xbar = 5.33
sample standard deviation, s = 1.85
sample size, n = 8
degrees of freedom, df = n - 1 = 7

Given CI level is 95%, hence α = 1 - 0.95 = 0.05
α/2 = 0.05/2 = 0.025, tc = t(α/2, df) = 2.365

ZCL = 2.365


b)

ME = tc * s/sqrt(n)
ME = 2.365 * 1.85/sqrt(8)
ME = 1.55